OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 4,7,v65 * v stands for sqrt A. yes B. no

OpenStudy (anonymous):

4^2+7^2 = 65 v65 ^2 = 65 therefore, yes

OpenStudy (eyust707):

\[a^2 + b^2 = c^2\] \[a=4\] \[b=7 \]